# MARVEL-40M+: Multi-Level Visual Elaboration for High-Fidelity Text-to-3D Content Creation

### v1

@@ Abstract
Generating high-fidelity 3D content from text prompts remains a significant challenge in computer vision due to the limited size, diversity, and annotation depth of the existing datasets. To address this, we introduce MARVEL-$40$M+, an extensive dataset with $40$ million text annotations for over $8.9$ million 3D assets aggregated from seven major 3D datasets. Our contribution is a novel multi-stage annotation pipeline that integrates open-source pretrained multi-view VLMs and LLMs to automatically produce multi-level descriptions, ranging from detailed ($150$-$200$ words) to concise semantic tags ($10$-$20$ words). This structure supports both fine-grained 3D reconstruction and rapid prototyping. Furthermore, we incorporate human metadata from source datasets into our annotation pipeline to add domain-specific information in our annotation and reduce VLM hallucinations. Additionally, we develop MARVEL-FX3D, a two-stage text-to-3D pipeline. We fine-tune Stable Diffusion with our annotations and use a pretrained image-to-3D network to generate 3D textured meshes within 15s. Extensive evaluations show that MARVEL-40M+ significantly outperforms existing datasets in annotation quality and linguistic diversity, achieving win rates of $72.41\%$ by GPT-4 and $73.40\%$ by human evaluators.
@@ 3.1 Multi-Stage Annotation Pipeline
Multi-View Rendering: We first generate $4$ multi-view images of resolution $512\times 512$ for each 3D model using Blender [1]. We rotate the camera around the object with azimuth angle, $\theta=\{\frac{\pi i}{2}\}_{i=1}^{i=4}$ and fixed elevation angle, $\phi=60$. The camera distance is set to $1.5$. The four images correspond to the front, back, left, and right sides of the 3D model. Unlike existing 3D captioning pipelines [53, 28], we focus solely on these standard viewpoints. This method aligns with recent studies [68, 79], which demonstrate that VLMs perform better on images from these viewpoints.
@@
Dense Description Generation: In this stage, InternVL2 [13, 15] processes the $4$ rendered multi-view images along with our metadata-augmented prompt to generate a dense description of the 3D models. This description contains several key requirements for fine-grained 3D model reconstruction: (1) structural decomposition with object identification and relative positions, (2) geometric properties, analyzing shape characteristics, symmetry axes, and proportional relationships, (3) surface characteristics, addressing texture and material properties and tactile qualities such as roughness and reflectivity; (4) chromatic analysis, mapping colors across primary objects and sub-components, including patterns and transitions (5) environmental context, capturing spatial relationships and its interaction with other elements.
@@
An example of multi-level visual elaboration is illustrated in Figure 2 (left), where Qwen $2.5$ [85] compresses texture information progressively from Level 1 to Level 4. By Level 5, the output shifts to a concise format, highlighted with colored words representing key semantic tags and core attributes. To assess the effectiveness of this hierarchical compression, we conduct an ablation study in Section 4.3 B, measuring how well semantic information is retained across all the levels.
@@
Figure 3: Qualitative Annotation Comparison: From top to bottom Cap3D [53], 3DTopia [28], Kabra [34], MARVEL (Level-4) annotations and GPT-4 [60] evaluation. MARVEL consistently provides the most comprehensive and precise annotations, capturing intricate details such as object names, color, structure, and specific attributes. Red is for wrong captions.
@@ 3.2 Caption Generation
Datasets: We aggregate 8.98.9~{}8.98.9M 3D assets from seven diverse sources [18, 17, 80, 74, 10, 16, 73, 20]. For human metadata injection, we use the name, tags and description from Objaverse 1.0 [18] (Sketchfab) and metadata from Objaverse-XL [18] (thingiverse and github). Samples from Objaverse-XL [17] containing the file extension .ply were excluded from the dataset. For the rest of the six datasets [10, 16, 20, 80, 73, 74], we use the class categories as metadata. Any samples that do not have renderable multi-views or lack textual information post-annotation are removed from the dataset. The final details of the dataset are provided in Table 1, with additional preparation information available in the supplementary materials.
@@ 3.3 MARVEL-FX3D Architecture
Image-to-3D Generation: In the second stage, the background is removed from the generated image using DIS [63]. The refined image is then processed by pretrained SF3D [7] to generate a high-quality textured mesh within 5s. Further details can be found in the supplementary.
@@ 4.1 Annotation Evaluation
Caption Accuracy: Table 3 shows the caption accuracy results, where MARVEL (Level 1) achieves the highest scores—$84.70\%$ in GPT-4 evaluation and $82.80\%$ in human evaluation—demonstrating superior consistency compared to other methods. Baseline datasets with shorter captions, like Kabra [34] and others [28, 53], tend to capture objects semantically (e.g., a statute,a man in a tuxedo) but lack detailed descriptions. Although longer captions increase the risk of errors, MARVEL (Level 1) maintains high accuracy with an average length of 170 words (34×34$\times$34 × that of Kabra [34], 10×10$\times$10 × that of Cap3D [53], and 5×5$\times$5 × that of 3D-Topia [28]), effectively balancing detail and correctness. As shown in Figure 3, MARVEL captures both domain-specific names and intricate features, exemplified by captions like The Monument of Dante Alighieri… flowing robes, holding a book. It stands on a decorated stone plinth and is made of smooth, polished marble…
@@ 4.2 Text-to-3D Generation
Implementation Details: We fine-tune SD $3.5$ [21, 3] using the Objaverse [18] dataset, which includes $798,759$ 3D assets, split into training, validation, and test sets in a $90:5:5$ ratio. Fine-tuning is conducted in half-precision for $5$ epochs with a batch size of $8$, using a single NVIDIA H100100100100 GPU, with LoRA [30] rank and alpha set to $4$. Further details are provided in the supplementary section.
Baselines: To evaluate MARVEL-FX3D’s performance in high-fidelity TT3D generation, we compare it with Shap-E [33], Dreamfusion [62], Luciddreamer [43], and HIFA [91]. We use the official implementations and pretrained models for Shap-E and Luciddreamer, training the latter for 3k steps. Dreamfusion and HIFA are trained using the open-source threestudio [23] implementation, with 10k and 24k steps, respectively, under default settings. Due to the slower optimization of Dreamfusion, HIFA, and Luciddreamer, we limit comparisons to 50 randomly selected samples from the Objaverse [18] test set. Instant3D [41] and Assetgen [71] are excluded due to unavailable code.
@@ 4.3 Ablation Study
A. Effect of Human Metadata on Annotation Quality: Human-generated metadata is vital in the MARVEL annotation pipeline, enriching text captions with domain-specific details. While quantitative analysis would require specialized expertise, we provide qualitative evidence of its impact. As shown in Figure 5, MARVEL accurately identifies specific details, such as three human footprints on a rocky surface, which both InternVL-2 [13, 15] and GPT-4 [60] miss, producing only generic descriptions. Similarly, MARVEL annotations include detailed identifiers like specific lunar craters, which are absent in outputs from InternVL-2 and GPT-4. This highlights how integrating human metadata enhances the context in annotations. Additional examples spanning other domains (e.g. biological elements, historical sites) are detailed in the supplementary material.
@@
Table 5: Ablation study results showing SCS across MARVEL-40M levels, illustrating strong semantic retention through Levels 1-4 and reduced detail at Level 5.
@@ 5 Limitation
Our analysis reveals some limitations in MARVEL annotation pipeline and MARVEL-FX3D. First, the underlying VLMs and LLMs exhibit inherent weaknesses in numerical precision [61, 9] and directional understanding [26] in complex scenes with multiple objects and occlusion. Second, InternVL-2 struggles with very thin objects, often misidentifying their side-views as different objects entirely. Finally, without metadata support, the caption accuracy becomes generic for complex 3D structures, particularly in scenes with fragmented geometries like architectural interiors. Additionally, MARVEL-FX3D sometimes generates flat 3D objects due to depth ambiguity in the input image. In supplementary section, some visual examples are provided. Despite these challenges, it is important to note that the strengths of our proposed pipeline and architecture remain significant, as they are both model-agnostic and adaptable to future enhancements.
@@ 6 Conclusion
In this work, we introduced MARVEL-$40$M+, the largest 3D captioning dataset to date, comprising over $40$ million high-quality text annotations for $8.9$ million 3D assets across seven major 3D datasets. Our primary contributions include a scalable, multi-stage annotation pipeline that combines open-source pretrained multi-view VLMs and LLMs with filtered human metadata to reduce hallucinations and introduce domain-specific information. Our pipeline produces five levels of annotations for diverse 3D modeling needs, from detailed reconstruction descriptions to rapid prototyping tags. Additionally, we introduce MARVEL-FX3D, a two-stage architecture that leverages fine-tuned Stable Diffusion on our dataset and pretrained Stable Fast 3D to generate high-quality, textured 3D meshes in just $15$ seconds. Through extensive experimentation, we demonstrated both MARVEL-$40$M+’s superior annotation quality and linguistic depth, and MARVEL-FX3D’s state-of-the-art performance in high fidelity text-to-3D generation. We believe that MARVEL-$40$M+ will serve as a foundational resource for future advancements in text-to-3D content creation, inspiring further research to address the current limitations and expand the dataset’s applications.
@@ 7 Acknowledgement
This work was in parts supported by the EU Horizon Europe Framework under grant agreement 101135724 (LUMINOUS).
@@
Figure 6: An example use case of MARVEL-FX3D, demonstrating how multiple prompts can be combined to create a detailed and complex 3D scene, with each prompt contributing specific elements such as characters, structures, and environmental details. Please zoom in for details.
@@ 8.1 Dataset Preparation
ShapeNet: For the ShapeNet dataset, which contains 51,209 samples, we use the ShapeNet taxonomy as its metadata (e.g., airplane, bowl, cap, clock, etc.).
Pix3D: For the Pix3Dhttp://pix3d.csail.mit.edu/[74] dataset, which contains 735 samples, we use the associated category tag as its metadata (e.g., bed, table, desk, chair, etc.).
@@ 8.2 Implementation Details
For human evaluations in our paper, we developed a Gradio app to compare our captions with those from baseline datasets, including Cap3D, 3DTopia, and Kabra, as well as to evaluate FX3D results against text-to-3D baselines. The evaluations were conducted by a panel of five human experts.

### v2

@@ Abstract
Generating high-fidelity 3D content from text prompts remains a significant challenge in computer vision due to the limited size, diversity, and annotation depth of the existing datasets. To address this, we introduce MARVEL-$40$M+, an extensive dataset with $40$ million text annotations for over $8.9$ million 3D assets aggregated from seven major 3D datasets. Our contribution is a novel multi-stage annotation pipeline that integrates open-source pretrained multi-view VLMs and LLMs to automatically produce multi-level descriptions, ranging from detailed ($150$-$200$ words) to concise semantic tags ($10$-$20$ words). This structure supports both fine-grained 3D reconstruction and rapid prototyping. Furthermore, we incorporate human metadata from source datasets into our annotation pipeline to add domain-specific information in our annotation and reduce VLM hallucinations. Additionally, we develop MARVEL-FX3D, a two-stage text-to-3D pipeline. We fine-tune Stable Diffusion with our annotations and use a pretrained image-to-3D network to generate 3D textured meshes within 15s. Extensive evaluations show that MARVEL-40M+ significantly outperforms existing datasets in annotation quality and linguistic diversity, achieving win rates of $72.41\%$ by GPT-4 and $73.40\%$ by human evaluators. Project page is available at https://sankalpsinha-cmos.github.io/MARVEL/.
@@ 3.1 Multi-Stage Annotation Pipeline
Multi-View Rendering: We first generate $4$ multi-view images of resolution $512\times 512$ for each 3D model using Blender [1]. We rotate the camera around the object with azimuth angle, $\theta=\{\frac{\pi i}{2}\}_{i=1}^{i=4}$ and fixed elevation angle, $\phi=30$. Models are scaled to a unit bounding box, centered at the origin, with the camera distance set as $1.5$ relative to this. The four images correspond to the front, back, left, and right sides of the 3D model. Unlike existing 3D captioning pipelines [53, 28], we focus solely on these standard viewpoints. This method aligns with recent studies [68, 79], which demonstrate that VLMs perform better on images from these viewpoints.
@@
Dense Description Generation: In this stage, InternVL2 [13, 15] takes as input the $4$ rendered multi-view images along with our metadata-augmented prompt to generate a dense description of the 3D models. This step avoids separate processing and view aggregation [53, 28] required by pipelines using models like LLaVA [45, 46] or BLIP-2 [39, 40]. The generated description contains several key requirements for fine-grained 3D model reconstruction: (1) structural decomposition with object identification and relative positions, (2) geometric properties, analyzing shape characteristics, symmetry axes, and proportional relationships, (3) surface characteristics, addressing texture and material properties and tactile qualities such as roughness and reflectivity; (4) chromatic analysis, mapping colors across primary objects and sub-components, including patterns and transitions (5) environmental context, capturing spatial relationships and its interaction with other elements.
@@
An example of multi-level visual elaboration is illustrated in Figure 2 (left), where Qwen $2.5$ [85] compresses texture information progressively from Level 1 to Level 4. By Level 5, the output shifts to a concise format with colored words representing key semantic tags and core attributes. To assess the effectiveness of this hierarchical compression, we conduct an ablation study in Section 4.3B, measuring the retention of semantic information across all levels.
@@
Figure 3: Qualitative Annotation Comparison: From top to bottom Cap3D [53], 3DTopia [28], Kabra [34], MARVEL (Level-4) annotations and GPT-4 [60] evaluation. MARVEL consistently provides the most comprehensive and precise annotations, capturing intricate details such as object names, color, structure, and specific attributes. Red is for wrong captions. Green shows important information.
@@ 3.2 Caption Generation
Datasets: We aggregate $\sim 8.9$M 3D assets from seven diverse sources [18, 17, 80, 74, 10, 16, 73, 20]. For human metadata injection, we use the name, tags and description from Objaverse 1.0 [18] (Sketchfab) and metadata from Objaverse-XL [18] (thingiverse and github). Samples from Objaverse-XL [17] containing the file extension .ply were excluded from the dataset. For the rest of the six datasets [10, 16, 20, 80, 73, 74], we use the class categories as metadata. Samples without any renderable multi-view images or contains zero-length final annotations are removed from the dataset. The final details of the dataset are provided in Table 1, with additional preparation information available in the supplementary materials.
@@ 3.3 MARVEL-FX3D Architecture
Image-to-3D Generation: In the second stage, the background is removed from the generated image using DIS [63]. The refined image is then passed to pretrained SF3D [7] to generate a high-quality textured mesh in 5s.
@@ 4.1 Annotation Evaluation
Caption Accuracy: Table 3 shows the caption accuracy results, where MARVEL (Level 1) achieves the highest scores—$84.70\%$ in GPT-4 evaluation and $82.80\%$ in human evaluation—demonstrating superior consistency compared to other methods. Baseline datasets with shorter captions, like Kabra [34] and others [28, 53], tend to capture objects semantically (e.g., a statute,a man in a tuxedo) but lack detailed descriptions. Although longer captions increase the risk of errors, MARVEL (Level 1) maintains high accuracy with an average length of 170 words (34×34$\times$34 × that of Kabra [34], 10×10$\times$10 × that of Cap3D [53], and 5×5$\times$5 × that of 3D-Topia [28]), effectively balancing detail and correctness. As shown in Figure 3, MARVEL captures both domain-specific names and intricate features, exemplified by captions like The Monument of Dante Alighieri… flowing robes, holding a book … smooth, polished marble…
@@ 4.2 Text-to-3D Generation
Implementation Details: We fine-tune SD $3.5$ [21, 3] using the Objaverse [18] dataset, which includes $798,759$ 3D assets, split into training, validation, and test sets in a $90:5:5$ ratio. Fine-tuning is conducted in half-precision for $5$ epochs with a batch size of $8$, using a single NVIDIA H100100100100 GPU, with LoRA [30] rank and alpha set to $4$.
Baselines: To assess MARVEL-FX3D’s performance in high-fidelity TT3D generation, we compare it with Shap-E [33], Dreamfusion [62], Luciddreamer [43], and HIFA [91]. Due to their slower optimization [62, 43, 91], we evaluate 50 random level 4 captions from the Objaverse [18] test set. Instant3D [41] and Assetgen [71] are excluded due to unavailable code. See supplementary for details.
@@
Cap3D vs MARVEL-40M+: To demonstrate the effectiveness of our dataset on TT3D generation, we conduct human evaluation on the performance of MARVEL-FX3D when SD $3.5$ is either pretrained or finetuned on MARVEL-40M+ and Cap3D [53] captions. The evaluation follows the same strategy and uses the same 50 test samples as before. Results on Table 5 confirm that MARVEL-FX3D trained on our captions consistently outperforms pretrained and Cap3D versions across all metrics, with a notable increase in prompt fidelity.
Dataset Text-to-Image [3] Geometric Visual Prompt Overall
Consistency Quality Fidelity
  Pretrained 2.51 ± 1.84 2.54 ± 1.78 2.58 ± 1.86 2.41 ± 1.69
Cap3D Finetuned 6.51 ± 1.48 6.53 ± 1.52 6.54 ± 1.68 6.43 ± 1.47
MARVEL Finetuned 7.20 ± 0.91 6.58 ± 0.86 7.71 ± 0.68 6.94 ± 0.71
Table 5: Quantitative evaluation of MARVEL-FX3D without and with fine-tuning on Cap3D and MARVEL captions, showing improved performance with MARVEL.
@@ 4.3 Ablation Study
A. Effect of Human Metadata on Annotation Quality: Human-generated metadata is vital in the MARVEL annotation pipeline, enriching text captions with domain-specific details. While quantitative analysis requires specialized expertise, we provide its qualitative evidence. As shown in Figure 5, MARVEL accurately identifies specific details, such as three human footprints on a rocky surface, which both InternVL-2 [13, 15] and GPT-4 [60] miss, producing only generic descriptions. Similarly, MARVEL captures identifiers like specific lunar craters, absent in InternVL-2 and GPT-4 outputs. This highlights how integrating human metadata enhances the context in annotations. Additional examples from other domains (e.g., biology, historical sites) are in the supplementary material.
@@
Table 6: MARVEL-40M+ illustrates strong semantic retention through Levels 1-4 and reduced detail at Level 5.
@@ 5 Limitation
Our analysis reveals some limitations in MARVEL annotation pipeline and MARVEL-FX3D. First, the underlying VLMs and LLMs struggle with numerical precision [61, 9] and directional understanding [26] in complex scenes with multiple objects and occlusion. Second, InternVL-2 misidentifies very thin objects, often treating side-views as separate entities. Finally, without metadata, captions become generic for complex 3D structures, especially in fragmented geometries like architectural interiors. In supplementary, some visual examples are provided. Additionally, MARVEL-FX3D sometimes generates flat 3D objects due to depth ambiguity in the input image. Despite these challenges, our pipeline remains model-agnostic and adaptable to future enhancements.
@@ 6 Conclusion
We present MARVEL-$40$M+, the largest 3D captioning dataset to date, comprising over $40$M+ high-quality text annotations for $8.9$ million 3D assets across seven major 3D datasets. Our primary contributions include a scalable, multi-stage annotation pipeline that combines open-source pretrained multi-view VLMs and LLMs with filtered human metadata to reduce hallucinations and introduce domain-specific information. Our pipeline produces five levels of annotations for diverse 3D modeling needs, from detailed descriptions to concise tags. Additionally, we introduce MARVEL-FX3D, a two-stage architecture that leverages fine-tuned Stable Diffusion on our dataset and pretrained Stable Fast 3D to generate high-quality, textured 3D meshes in just $15$s. Through extensive experimentation, we demonstrated both MARVEL-$40$M+’s superior annotation quality and linguistic depth, and MARVEL-FX3D’s improved performance in high fidelity TT3D generation. We believe MARVEL-$40$M+ will be a foundational resource for advancing TT3D content creation, inspiring further research to address current limitations and expand its applications.
@@ 7 Acknowledgement
This work was co-funded by the European Union under Horizon Europe, grant number $101135724$, project LUMINOUS. However, the views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible. Human evaluation for this project was primarily funded by BITS Pilani Hyderabad’s NFSG Grant (Reference N4/24/1033).
@@
Figure 6: An example use case of MARVEL-FX3D, demonstrating how multiple prompts can be combined to create a detailed and complex 3D scene, with each prompt contributing specific elements such as characters, structures, and environmental details (Zoom in for details).
@@ 8.1 Dataset Preparation
ShapeNet: For the ShapeNet dataset, which contains 52,472 samples, we use the ShapeNet taxonomy as its metadata (e.g., airplane, bowl, cap, clock, etc.).
Pix3D: For the Pix3Dhttp://pix3d.csail.mit.edu/[74] dataset, which contains 374 samples, we use the associated category tag as its metadata (e.g., bed, table, desk, chair, etc.).
@@ 8.2 Implementation Details
For human evaluations in our paper, we developed a Gradio app to compare our captions with those from baseline datasets, including Cap3D, 3DTopia, and Kabra, as well as to evaluate FX3D results against text-to-3D baselines. The evaluations were conducted by a panel of 5 human experts.
8.3 Compute and GPU Hours
MARVEL’s annotation pipeline utilizes one NVIDIA H100-80GB GPU, one RTX-4090 GPU, and one RTX-A6000 GPU, achieving a throughput of approximately 24,000 samples per day. Annotating the entire Objaverse dataset (800,000 samples) would thus require about 33 days, incurring an estimated total computational cost of approximately \$2,700–\$3,000, based on publicly available GPU pricing‡‡‡‡‡‡https://tinyurl.com/gpu-usage-pricing (Original).
For comparison, sequential human annotation has a considerably lower throughput (1,400 samples/day) and higher cost (\$87.18 per 1,000 annotations), resulting in approximately 572 days (about 1.57 years) and a total cost of roughly \$69,744 for annotating the complete Objaverse dataset. In contrast, the automated Cap3D pipeline—leveraging BLIP2, CLIP, and GPT-4 models on cloud-hosted NVIDIA A40 GPUs—achieves significantly higher throughput (65,000 samples/day) at a lower cost (\$8.35 per 1k annotations), requiring only about 13 days and totaling approximately \$6,680 for the entire dataset [53].
Our pipeline annotates the Objaverse dataset at approximately half the total cost of Cap3D, although with a lower throughput (33 days vs. Cap3D’s 13 days). Both automated methods substantially outperform sequential human annotation in terms of speed and cost. Importantly, our pipeline delivers annotations of significantly higher quality compared to Cap3D, making it particularly advantageous when balancing annotation quality and cost efficiency. All comparisons assume sequential (non-parallelized) processing; parallelization would further reduce annotation time for all methods.
Method
Throughput
(samples/day)
Total Days
(800k samples)
Cost per 1k
annotations
Total Cost
(800k samples)
Human 1,400 572 \$87.18 \$69,744
Cap3D 65,000 13 \$8.35 \$6,680
MARVEL 24,000 33 \$3.38–\$3.75 \$2,700–\$3,000
Table 7: Comparison of annotation pipelines based on throughput, annotation time, and cost for annotating the Objaverse dataset (800k samples). All estimates assume sequential annotation without parallelization.
@@ 9.3 More Multi-Level Examples
9.4 More on Simple and Textureless Models
Our annotation pipeline is robust and adaptable, effectively handling both simple and texture-less models by dynamically adjusting its descriptive verbosity. As illustrated in Figure 22, the pipeline generates concise yet accurate annotations for texture-less models—such as the smooth, monochromatic humanoid figure and the symmetrical lemon head with sunglasses—by emphasizing their geometric precision and structural symmetry. Additionally, for simpler models like the low-poly tree with geometric leaves and the realistically textured orange, the pipeline adeptly captures essential details, highlighting subtle irregularities and primary shape characteristics. This flexibility ensures consistent annotation quality across diverse modeling scenarios.
9.5 Need for Multi-Level-Structure
Hierarchical structures are essential in AI. Examples include multi-resolution models in computer vision, such as feature pyramids, and hierarchical embeddings in NLP, such as document summarization. In 3D modeling, ShapeNet uses a hierarchical taxonomy [10] to improve adaptability across tasks. MARVEL similarly adopts a hierarchical design, ensuring task-specific granularity. By using a predefined hierarchy, MARVEL eliminates the need for repeated prompting. This reduces latency, computational costs, and inconsistencies associated with dynamically adjusting verbosity. Such dynamic adjustments would require multiple inference steps and additional processing, making them impractical for large-scale pipelines—even for future LLMs/VLMs. Additionally, dynamic generation introduces risks like semantic drift, information loss, and verbosity imbalance, decreasing annotation reliability. Our ablation study (Section 4.3B and Table 5) confirms that MARVEL’s structured verbosity effectively maintains essential details. It optimizes verbosity levels according to task requirements, as validated by cosine similarity and compression ratio.
@@ 9.6 Failure Cases
9.7 More on MTLD Scores
The MTLD (Measure of Textual Lexical Diversity) algorithm quantifies vocabulary diversity by segmenting a text whenever the Type-Token Ratio (TTR)—the ratio of unique words to total words—drops below a fixed threshold (commonly 0.72). It processes the text both forwards and in reverse to reduce positional bias, and calculates the final MTLD score as the total number of words divided by the number of segments (called factors). A low MTLD score indicates a repetitive vocabulary and low lexical diversity, while a high score reflects a rich and varied vocabulary. For instance, the repetitive string "hello hello hello hello hello hello" results in a low MTLD score of approximately 2.02, due to the lack of word variation. In contrast, the diverse sentence "the quick brown fox jumps over the lazy dog" yields a high MTLD score of around 22.68, as it contains many unique words. The pseudo-code for the algorithm is given in Algorithm 1 as seen in [56].
Algorithm 1 MTLD Score [56]
1:function MTLD(text, min = 10)
2:       $forward\leftarrow\textsc{MTLDProcess}(text,min)$
3:       $reverse\leftarrow\textsc{MTLDProcess}(Reverse\_text,min)$
4:       return $(forward+reverse)/2$
5:end function
6:function MTLDProcess(text, min)
7:       $factor\leftarrow 0$
8:       $factor\_lengths\leftarrow 0$
9:       $start\leftarrow 0$
10:       for $x\leftarrow 0$ to $length(text)-1$ do
11:             $segment\leftarrow text[start:x+1]$
12:             if $x+1=length(text)$ then
13:                    $partial\leftarrow\frac{1-\textsc{TTR}(segment)}{1-0.72}$
14:                    $factor\leftarrow factor+partial$
15:                    $factor\_lengths\leftarrow factor\_lengths+length(segment)$
16:             else if $\textsc{TTR}(segment)<0.72$ and $length(segment)\geq min$ then
17:                    $factor\leftarrow factor+1$
18:                    $factor\_lengths\leftarrow factor\_lengths+length(segment)$
19:                    $start\leftarrow x+1$
20:             end if
21:       end for
22:       return $\frac{factor\_lengths}{factor}$
23:end function